July 1, 2015 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, July 1, 2015
11:00 am, EDT
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2015 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2015/ .


  1. Announcements
  2. Subdetector Reports
  3. Detector Matching
  4. Simulations
  5. Calibration Processing
    • Run quality criteria
  6. AOB

Action Items

  • Produce updated calibration plots and performance numbers, preferably within a week.
  • Polish calibration routines, including testing of results, documentation (GlueX note).
  • Update simulation parameters.


Attending: Sean (NU); Curtis, Mike S., Will M., Naomi (CMU); Simon, Adesh, Beni, Matt, Kei, Mahmoud, Justin, Eric, Alex B., Fernando, Luke, Mark I., Paul, Elton, Mark D.


  • The JLab S&T review will be at the end of the month, so updated plots showing detector calibration and performance and updated numbers for resolutions and efficiencies are needed. Matt suggested producing one slide per subdetector, including PDF files for each figure and some rough text summarizing the overall status and contents of the figures. This information can be adapted by the various speakers as needed, and is needed by mid-next week.
  • The next offline monitoring run is planned for this coming Monday, July 6.


  • Simon showed results of a study using tracks matched to FCAL showers to extract relative timing offsets for FCAL blocks. pion/proton separation is seen.
    • Matt pointed out that there were timing offsets extracted from LED data already in the CCDB, but they were overwritten during recent timing updates. Adesh and Simon will compare these two results.
    • Simon mentioned that these studies were motivated by noticing that sometimes a cluster will be reconstructed by two different hits in the same block with different times. Matt agreed that this should be studied.
  • Matt and Adesh are investigating the non-linear energy-dependent corrections and plan to have updated constants installed in the CCDB by the upcoming offline monitoring run, along with update gain calibrations.


  • Elton reported that George is converging on some results for his studies of effective velocities.
    • Another output of this procedure is relative time offsets between the upstream and downstream channels. The procedure uses ADC times. It was agreed that calibrating absolute time offsets for the individual channels was a next step to take.
    • Curtis asked if TDC times were being analyzed, and Elton responded that the first step in using these is to determine the timewalk corrections for each channel. Noemi from Regina is working on these, but does not have results from data yet.


  • Simon is updating his CDC alignment. He hasn't sent his results to Mike S. yet, but promises to do soon.
  • Mike S. is checking his CDC alignment results using Millipede with simulated data, and has removed several features from the results (e.g. oscillations in a ring). He is getting reasonable input/output agreement of the wire offsets, within 50um.
  • Lubomir was not here, so no FDC news.


  • Beni is redoing his calibration results, since Simon pointed out a small problem with his code. It was suggested that he use the common library functions for converting from TDC digitized time to uncalibrated time in physical units.
  • Now that the counting house is back online, Beni has resumed his studies into the fADC data.

Start Counter

Eric listed a number of projects he is working on:

  • Adding in the use of the attenuation factors to the reconstruction code
  • Seeing if the propagation time corrections derived from bench measurements or beam data best minimize the difference between the Start Counter and RF times, in order to decide which corrections to use.
  • Obtained new timewalk corrections. These are very similar to the timewalk corrections previously obtained, but are obtained with a new fitting procedure that produces more regular, stable results.


  • Alex B. is looking into using the tagger/PS energies to obtain efficiencies from the data. An updated PS magnet map is needed for the most accurate results. No new news from Alex Somov on this front.

Other Issues

  • Matt asked if we should worry about calibrating both fall and spring data. Sean said that at a previous calibration meeting, it had been agreed to focus calibrations on the spring data, since the quality of the data is much better. After some general discussion, it was agreed to drop the Fall 2014 data from the regular offline monitoring runs. This data can be re-processed on demand.
  • Fernando reported that he has continued studies of pedestals on the fADC250. He will upload a document describing the results to the DocDB soon, but the main result is that 16 is the minimum number of samples needed for a good determination of the pedestal.
  • Related to this, Beni mentioned that he looked at data taken with the fADC250/125, and found that while calculating the pedestal using 4 samples, there was a difference between using floating point and integer arithmetic (as is used on the FPGAs), but not when using 16 samples.

Detector Matching

Sean reviewed some monitoring plots from the last monitoring run. There was much discussion Some comments:

  • This was a low current beam run, so there could be substantial contributions from cosmic rays in the CDC area (either in-time cosmic rays or ones that triggered the BCAL).
  • The track/start counter matching could be improved in the nose region by, e.g., loosening the delta-phi cut in a z-dependent fashion.
  • The TOF still does not have channel-by-channel time offsets. The geometry probably needs to be updated - Simon will look into this.
  • The FCAL matching calculations shown probably depend strongly on the TOF matching performance.
  • The tracks used in the inclusive pi/p distributions shown are selected using PID FOM cuts, and don't show general features of the tracking performance. Sean will look into this in more detail.


Mark I. generated a new software release, and started running test jobs for the various beam conditions on Monday. Paul found a problem with how the TOF information was being written out, and the software is being updated and the test jobs redone.

Calibration Processing

Sean briefly discussed the ongoing work on calibration processing. He is working on incorporating various calibrations into this scheme. He highlighted that in order to generate semi-automatic calibrations, a general goal for completed calibration procedures are:

  • Plugins should be stored in standard location in sim-recon tree
  • ROOT and other scripts stored in standard location in SVN
  • Output to files should be defined
  • QA routines should be developed